# Ravnest: Decentralized Asynchronous Training on Heterogeneous Devices

### v1

@@ 1 Introduction
Towards distributing the training process across multiple systems, Data Parallelism has been a very effective mechanism in which each system is responsible for a portion of the training data batch [49, 14]. In Synchronous Data Parallelism, the workers have to wait for gradients to be computed across all peers so that parameter updates can take place and the next data batch can be processed. However, the presence of workers that are much slower than others will significantly impede efficiency. To overcome this, Asynchronous mechanisms were introduced where peers don’t have to wait for each other to make parameter updates [46]. Though Asynchronous mechanisms may give rise to updates with stale gradients, they have been observed to converge efficiently under certain constraints [10]. In distributed training setups, communication-efficient All-Reduce mechanisms, such as the widely-used Nvidia Collective Communications Library (NCCL) [26, 18, 37], are preferred for data transmission between peers. This approach is favored over a common parameter server, which can become a bottleneck when the number of communicating peers is significantly large.
@@
This paper details an Asynchronous Parallel training approach that utilizes both Data and Model Parallelism, allowing weaker systems to participate in distributed training while minimizing the idle time bubble that affects the efficiency of Synchronous Model Parallelism. The proposed method eliminates the need for workers to host entire models and mitigates the reduction in efficiency due to stragglers by grouping workers into clusters within which Zero-Bubble Asynchronous Model Parallelism is executed. Periodic parameter synchronisation across clusters is conducted via a communication efficient All-Reduce mechanism. Theoretical Analysis and derivations for convergence rate in terms of the number of updates ($K$) show consistencies with Local SGD. We also discuss linear speedup with respect to the number of participating clusters and constraints on the delay parameter under which this can be achieved. This mechanism can efficiently train large state of the art models while significantly reducing hardware requirements and training costs.
@@
  • Ravnest, a novel asynchronous parallel training approach that amalgamates the best features of data and model parallelism to facilitate the distributed training of complex deep learning models over large datasets on clusters of heterogeneous consumer grade PCs over the internet, is introduced. With scalability and performance in mind, it aims to empower researchers and ML practitioners to develop and deploy deep learning models with ease, opening up new possibilities for groundbreaking research and real-world applications.
@@
  • The paper also explores how the cluster formation procedure can be framed as a constrained optimization problem and solved using meta heuristic approaches like Genetic Algorithm. This ensures that clusters with comparable data processing speeds are formed which leads to favourable conditions for data parallelism across clusters.
@@ 2 Related Works
2.1 Data Parallelism
Recent years have witnessed a staggeringly exponential growth in the size of datasets that are being used to train state of the art deep learning models. Storing such datasets spanning from terrabytes to petabytes in size poses a challenge in training models. This exact problem paved way to the creation of data parallelism, a technique that shards large datasets across multiple nodes, each of which host the entire model [35, 40]. Frameworks such like PyTorch Distributed Data Parallel [21] achieve this by first consigning copies of the entire model to individual accelerator nodes in high performance clusters, followed by sharding the dataset and iteratively dispatching chunks to corresponding machines. Copies of the complete model, hosted on each machine parallely performs forward and backward passes to derive the gradients of objective function on their corresponding shard of the dataset. Following this all gradients from individual machines are aggregated together during backpropagation to perform an optimization step. This approach presents some advantages such as full utilization of compute resources, faster training times [52], scalable infrastructures and support for synchronous and asynchronous updates based on the problem requirements. Despite its advantages, it is evident that on weaker systems with limited memory resources, hosting the entire may not necessarily be feasible.
2.2 Model Parallelism
With the rise of large models having complex architectures spanning terrabytes of data, there has been an inherent requirement to efficiently store and train them. Owing to constrained memory limits, the conventional training paradigms on centralized systems fail at this scale. Active research in recent times led to the development of a new paradigm of decentralized training called Model Parallelism, which can be used to train such massive models across a pool of devices wherein each device only holds a chunk of the main model (submodel) that it is capable of storing [47, 25, 15]. In this approach, we define a pipeline architecture of GPU enabled devices connected across a network. Each device hosts a submodel which takes in a batch of data, performs a collection of operations on this data to generate an output. This output is then fed as input to the subsequent device in the pipeline where another set of operations are performed and the cycle is repeated throughout the pipeline in a sequential manner. The network traffic and compute demand of the aforementioned operations are 2 key components that play a significant role in scalability of this approach. The sequential nature of standard model parallelism also poses a huge problem as it leads to under-utilization of the available compute resources. This can be observed in Figure 1a where the idle bubble represents a period of timesteps at which a majority of the devices have to stay idle while only one among them is busy performing a forward/backward pass.
@@ 2.2 Model Parallelism
To mitigate this issue, pipeline parallelism was developed as an improvement which splits the mini-batches into even smaller micro-batches and pipelines the computation over these micro-batches over multiple devices at the same time (as shown in Figure 1b), thereby reducing the idle time bubble and resulting in almost linear speedup [17]. [54] improves efficiency by overlapping an optimized cross-mesh resharding communication technique with the computations on the workers thereby reducing latency. As for an asynchronous setting, [16] systematically partitions the DNN layers to keep all devices active while maintaining a balance between communication across devices and computation on each individual device. While this significantly reduces the overall bubble period by scheduling the forward and backward passes in a round robin fashion, multiple machines remain idle during the startup state. [45] also successfully reduces the idle time bubble, but requires recomputation of forward passes as gradients are received in intermediate nodes to smoothen the effects of stale gradients.
@@ 2.3 Asynchronous Stochastic Gradient Descent
Initially, data parallelism based training mechanisms were facilitated by Synchronous Stochastic Gradient Descent (SSGD), which periodically synchronized the gradients of all the workers before each parameter update. This however, meant that the workers had to wait for each other before performing their updates, leading to deceleration in training and increased idle times created by slower workers [4].
On the other hand, Asynchronous Stochastic Gradient Descent (ASGD) allows workers to continue their individual training by allowing updates with stale gradients, reducing the extent to which workers have to wait for each other. Various improvements on vanilla ASGD, such as modified parameter or gradient averaging mechanisms, approximation functions to compensate for staleness and constraining the degree of staleness allowed in update procedures show that ASGD can match Synchronous methods on non-convex optimization problems. For instance, [53] leverages Taylor expansion of the gradient function and an efficient approximation to the Hessian matrix of the loss function to make the optimization behavior of ASGD closer to that of sequential SGD. [16, 45] show that Asynchronous training mechanisms with Model Parallelism converge efficiently while reducing system idle periods while mitigating the issues caused by straggling peers.
Distributed training traditionally utilized a Parameter Server which was responsible for gradient aggregation and weight updates [20]. The workers would frequently communicate with a dedicated Parameter Server to send their gradients and retrieve updated weights to progress with their individual training processes [30, 50, 51]. However, bottleneck and scalability issues arise since the parameter server has to regularly send and receive large payloads of sizes nearing the entire model, from all workers [11].
In Decentralized implementations of Stochastic Gradient Descent, each worker updates its local model by communicating with its peers. All-Reduce mechanisms are widely utilized for efficient averaging between all the participating peers [7] and are often preferred over gossip based methods due to their decrease in efficiency with an increase in the number of peers [33, 39]. Many implementations also use reduction mechanisms that limit the number of neighbours that a worker needs to communicate with, sometimes spanning multiple rounds of mixing, in order to reduce network traffic [6, 31, 38].
@@ 3.5.2 Cluster Formation Methodology
where $r_{i}$ and $b_{i}$ represents the RAM and steady bandwidth of the $i^{\text{th}}$ compute node. Let’s say we form $Q$ optimal clusters. Our aim is to find a new set $C=\left[c_{i}\right]$ where $c_{i}$ can take discrete values in $\left[1,2,...,Q\right]$. In other words, every $(r_{i},b_{i})$ is mapped to a corresponding discrete valued $c_{i}$ which represents the ID of the cluster that $(r_{i},b_{i})$ belongs to. Note that $P$ and $C$ have the same number of elements.
Since we need to effectively fit the entire model of size $M$ on each cluster, the first constraint in our optimization problem is that the sum of all peer’s RAM values need to be greater than $M$.
Across all clusters in a training session, the sum of individual times to transfer data taken by each node in a cluster should be comparable. This is equivalent to saying that the maximum difference in sums of data transfer times across any 2 clusters needs to be minimized.This forms our second constraint.
Since this is an NP-hard problem(meaning there’s no polynomial time solution), a heuristic/greedy approach might be better suited for finding optimal solutions.
This optimization problem can be solved using a Meta-Heuristic Genetic Algorithm [13]. We initialize a random population of individual chromosomes, iteratively evolve them across generations while introducing mutations and select the best performing ones based on a fitness function to create new successor generations [19].
The aforementioned set $C$ serves as a representation for an individual chromosome, which are assigned random values initially. A collection of such chromosomes are spawned to form a population, which are together evolved across generations. A penalty based fitness function (that uses the problem constraints) is then used to evaluate all individuals in a population out of which, the best ones are selected for crossover, resulting in offsprings that make up the next generation’s population. The individual with the best fitness value observed is tracked across all generations and returned.
Since each cluster formed will typically contain a mix of nodes with varying RAM values, the main model parameters will undergo a proportionate non-uniform split such that each peer can accomodate the largest chunk possible.

### v2

@@ 1 Introduction
Towards distributing the training process across multiple systems, Data Parallelism has been a very effective mechanism in which each system is responsible for a portion of the training data batch [54, 15]. In Synchronous Data Parallelism, the workers have to wait for gradients to be computed across all peers so that parameter updates can take place and the next data batch can be processed. However, the presence of workers that are much slower than others will significantly impede efficiency. To overcome this, asynchronous mechanisms were introduced where peers don’t have to wait for each other to make parameter updates [51]. Though asynchronous mechanisms may give rise to updates with stale gradients, they have been observed to converge efficiently under certain constraints [11]. In distributed training setups, communication-efficient All-Reduce mechanisms, such as the widely-used Nvidia Collective Communications Library (NCCL) [29, 41], are preferred for data transmission between peers. This approach is favored over a common parameter server, which can become a bottleneck when the number of communicating peers is significantly large.
@@
This paper details an asynchronous decentralized parallel training approach that utilizes both data and model parallelism, allowing weaker systems to participate in distributed training while minimizing idle time bubble. The proposed method eliminates the need for workers to host entire models and mitigates the reduction in efficiency due to stragglers by grouping workers into clusters within which Zero-Bubble Asynchronous Model Parallelism is executed. Periodic parameter synchronisation across clusters is conducted via a communication efficient Parallel Multi-Ring All-Reduce mechanism. Theoretical analysis and derivations for convergence rate in terms of the number of updates ($K$) show consistencies with local SGD. We also discuss linear speedup with respect to the number of participating clusters and constraints on the delay parameter under which this can be achieved. This mechanism can efficiently train large state of the art models while significantly reducing hardware requirements and training costs, opening up new possibilities for groundbreaking research and real-world applications.
@@
  • Ravnest, a novel asynchronous parallel training approach that amalgamates the best features of data and model parallelism to facilitate the distributed training of complex deep learning models over large datasets on clusters of heterogeneous consumer grade PCs over the internet, is introduced.
@@
  • Our experiments show significant memory reductions and competitive convergence rates with conventional synchronous training, complementing the theoretical analysis.
@@ 2 Related Works
2.1 Data and Model Parallelism
Due to exponential growth in dataset and model sizes, the need for parallelization and memory usage reduction led to the development of data parallelism. This technique involves using multiple systems, each hosting copies of the entire model, to process portions of a larger data batch in parallel followed by gradient aggregation and parameter updates [38, 44, 23]. This approach presents some advantages such as smaller batch sizes per system, full utilization of compute resources, faster training times [57], scalable infrastructures and support for synchronous and asynchronous updates based on the problem requirements. However, it is evident that on weaker systems with limited memory resources, hosting the entire model may not necessarily be feasible.
@@ 2.1 Data and Model Parallelism
Recent research has introduced model parallelism, a decentralized training paradigm enabling large model training across multiple devices, with each device holding only a portion of the main model it can store [52, 28, 16]. A pipeline of GPU devices is created with each device hosting a submodel processing data batches and passing results to the next device for further operations. Network traffic and compute demand are key factors for scalability. However, this sequential nature leads to under-utilization of compute resources since there occurs an idle bubble of timesteps as can be observed in Figure 1a. To address the issue, pipeline parallelism divides mini-batches into smaller micro-batches, processing them simultaneously across multiple devices (as shown in Figure 1b), thereby reducing the idle time bubble and resulting in almost linear speedup [19]. Pipeline parallelism inherently requires the participating devices to regularly communicate micro-batch results with each other. Recent works have improved communication efficiency by overlapping an optimized cross-mesh resharding technique with the computations on the workers, thereby reducing latency [59]. Asynchronously performing the forward and backward jobs at each pipeline stage while maintaining a balance between communication and computation can significantly reduce communication overhead and idle bubble time [17]. This has been achieved by scheduling the forward and backward passes in a round robin fashion. Significant reduction in idle time bubble has been successfully achieved in CNN architectures by considering varying degrees of staleness for each layer [50].
@@ 2.2 Asynchronous Stochastic Gradient Descent
Initially, data parallelism relied on Synchronous Stochastic Gradient Descent (SSGD), which periodically synchronized all workers’ gradients before each parameter update. However, this caused delays as workers waited for each other, slowing training and increasing idle time [4]. Asynchronous Stochastic Gradient Descent (ASGD) lets workers train independently by accepting updates with outdated gradients while minimizing wait times caused by straggling peers. Enhanced versions include adjusted averaging mechanisms and approximation functions to handle staleness, enabling ASGD to rival synchronous methods in non-convex optimization problems [58, 17, 50]. Distributed training traditionally relied on workers communicating with a dedicated parameter server for gradient aggregation and weight updates [22, 33, 55, 56]. However, bottleneck and scalability issues arise since the parameter server has to transmit large model-sized payloads to all workers regularly [12]. Workers update their local models by communicating with peers, often using all-reduce mechanisms for efficient averaging, which are preferred over gossip-based methods due to their efficiency with more peers [7, 36, 43]. Reduction mechanisms are also employed to limit worker communication, spanning multiple rounds to minimize network traffic [6, 34, 42].
@@ 3.5.2 Cluster Formation Methodology
where $r_{i}$ and $b_{i}$ represents the RAM and steady bandwidth of the $i^{\text{th}}$ compute node. Let’s say we form $Q$ optimal clusters. Our aim is to find a new set $C=\left[c_{i}\right]$ where $c_{i}$ can take discrete values in $\left[1,2,...,Q\right]$. In other words, every $(r_{i},b_{i})$ is mapped to a corresponding discrete valued $c_{i}$ which represents the ID of the cluster that $(r_{i},b_{i})$ belongs to. Note that $P$ and $C$ have the same number of elements. Since we need to effectively fit the entire model of size $M$ on each cluster, the first constraint in our optimization problem is that the sum of all peer’s RAM values need to be greater than $M$. Across all clusters in a training session, the sum of individual times to transfer data taken by each node in a cluster should be comparable. This is equivalent to saying that the maximum difference in sums of data transfer times across any 2 clusters needs to be minimized.This forms our second constraint.
This is an NP-hard problem(meaning there’s no polynomial time solution), a heuristic/greedy approach might be better suited for finding optimal solutions. This optimization problem can be solved using a Meta-Heuristic Genetic Algorithm [14]. We initialize a random population of individual chromosomes, iteratively evolve them across generations while introducing mutations and select the best performing ones based on a fitness function to create new successor generations [20].
The aforementioned set $C$ serves as a representation for an individual chromosome, which are assigned random values initially. A collection of such chromosomes are spawned to form a population, which are together evolved across generations. A penalty based fitness function (that uses the problem constraints) is then used to evaluate all individuals in a population out of which, the best ones are selected for crossover, resulting in offsprings that make up the next generation’s population. The individual with the best fitness value observed is tracked across all generations and returned. Since each cluster formed will typically contain a mix of nodes with varying RAM values, the main model parameters will undergo a proportionate non-uniform split such that each peer can accomodate the largest chunk possible.
@@
4 Experiments
This section elucidates upon a series of experimental training sessions conducted to evaluate the performance and efficiency of our proposed methodology across both single and multi-cluster setups with comparisons to baseline synchronous training.
4.1 Single Cluster Experiments with Homogeneous Nodes
We trained a ResNet-50 network [18] on the Tiny ImageNet dataset [48] for 50 epochs with a batch size of 100. SGD was utilized with a momentum of 0.9, weight decay of 0.0005 and learning rate of 0.01 which was reduced by a factor of 10 every 30 epochs. Images were pre-processed by scaling to 224x224x3, then normalized to [0,1], with a mean of [0.485, 0.456, 0.406] and standard deviation of [0.229, 0.224, 0.225]. We also trained an Inception-V3 model [40] on the CIFAR-10 dataset [21] for 50 epochs with a batch size of 64 and SGD with a momentum of 0.9, weight decay of 0.0005 and learning rate of 0.01, reducing by a factor of 10 every 30 epochs. Images were pre-processed by normalization (mean: [0.4914, 0.4822, 0.4465], standard deviation: [0.2023, 0.1994, 0.2010]). Augmentation techniques including random flipping, cropping and padding were applied for both datasets. While these settings provide a robust starting point, it is important to note that further fine-tuning of hyperparameters could potentially lead to improved performance.
(a) ResNet-50 model trained on Tiny ImageNet
(b) Inception-V3 model trained on CIFAR-10
Figure 5: Comparison of validation accuracies over epochs
For both models, conventional synchronous training was performed as a baseline, followed by decentralized training using proposed approach across 2 and 3 homogeneous nodes (within a single cluster), all hosted on Nvidia A10G GPUs with 24 GB VRAM. From Figures 5a and 5b, it is evident that as the number of nodes within the single cluster grows, initial validation accuracy registers lower values compared to scenarios with fewer nodes. This can be ascribed to the accumulation of more stagnant gradients as the size of the cluster increases. However, as training advances, the rates of convergence are at par with those of the baselines. We also observed that the average memory usage per node reduced by around 56% (3 nodes case) and by 42% (2 nodes case) for both models when compared to conventional training on a single GPU instance with the same batch sizes.
4.2 Multi-Cluster Experiments with Heterogeneous Nodes
We evaluated a multi-cluster setup using the proposed method on the BERT model following the BERT base (uncased) architecture [9]. Pretraining was done on Masked LM (MLM) and Next Sentence Prediction (NSP) tasks using the LAMB optimizer [54] with a learning rate of 0.00176 and weight decay of 0.01, on the WikiText-103 dataset [27]. A linear scheduler with 5000 warmup steps was also utilized. The sum of the losses for these 2 tasks was used as the convergence metric. The multi-cluster setup consisted of 10 nodes across A10G, V100 and T4 GPUs hosted on cloud instances. They were grouped into 4 clusters by the cluster formation genetic algorithm in order to ensure that the total compute times were similar across all clusters.
Figure 6: BERT pre-training on multi-cluster setup
Each cluster consisted of either 2 or 3 nodes, allowing for varying degrees of staleness, and gradients were accumulated over 32 batches of 8 samples before performing updates. Parameter averaging was done whenever each cluster completed 1400 update iterations. Conventional synchronous baseline training was done with the same hyperparameters. Figure 6 shows that in 11797 update iterations, the proposed algorithm achieves a learning curve comparable to that of the baseline. Training for more iterations or increasing the number of clusters can lead to even better results. Additionally, we noted that the average GPU memory usage across all nodes was approximately 67% less than that seen during training on a single GPU instance with the same batch size of 8.